Using Thevenins method to find Vx in the circuit below: -120 80 60 40 220 v z o 100 520A
Answer #1

Giren circuit is ojan o j68 Hun roso . mt, 35 520 4023år Thevenins equivalent we have to find for this circuitor jor 29 Boston lon sol v Sloxio - So Loo 8 jan eerma SOL) Hous using when super position theorem only holio is active -Lummis active when -2; only solo varen - O 50com Viño Bozon (-2;) 8+63 +10-2) 2 50 Loox -25 18+45 = -1.176 - 5.29; = 5.419 L-102.کر یہ دردز -)۱) (۱۰ +چند 3 (زد. ۱) (زی+ بی) = دند.) (زه +18) - زد. زه + و - عنه - ز36 = زو3- ها ز++18 زحام و ده. 2 - . ه = beVaa Van x . 3 3+ Req - 48.09 1-32.723 - 370.21~2.05j > 48.09 L-32.724 3.21~2.05) - 37.88 -0.103 i = 37.894-0.135°